# The Basic Reproduction Number for Petri Net Models: A Next-Generation Matrix Approach

### v2

@@ Abstract
The basic reproduction number ($R_{0}$) is an epidemiological metric that represents the average number of new infections caused by a single infectious individual in a completely susceptible population. The methodology for calculating this metric is well-defined for numerous model types, including, most prominently, Ordinary Differential Equations (ODEs). The basic reproduction number is used in disease modeling to predict the potential of an outbreak and the transmissibility of a disease, as well as by governments to inform public health interventions and resource allocation for controlling the spread of diseases. A Petri net (PN) is a directed bipartite graph where places, transitions, arcs, and the firing of the arcs determine the dynamic behavior of the system. Petri net models have been an increasingly used tool within the epidemiology community. However, a generalized method for calculating $R_{0}$ directly from PN models has not been established. Thus, in this paper, we present a general method for calculating $R_{0}$ for Petri nets. Additionally, we show how a computational method implementing the next-generation algorithm in ODE models can also be applied to Petri net models. We also provide multiple examples of how to use this approach to calculate $R_{0}$ for various SIR-type Petri net models.
keywords: Petri Nets, Reproduction Number, SIR model, Disease Spread, Next-Generation Method, Ordinary Differential Equations, Epidemiology, Modeling
@@ 1 Introduction
In the context of epidemiology, the basic reproduction number ($R_{0}$) refers to the expected number of new cases that one infected individual is likely to cause in a population where the entire population is susceptible [1]. As such, $R_{0}$ is a fundamental concept in mathematical epidemiology for all communicable diseases when it comes to Ordinary Differential Equation (ODE) based models. Intuitively, $R_{0}$, represents the expected number of secondary infections caused by a single infected individual in a fully susceptible population [1, 2]. An accurate calculation of $R_{0}$ is therefore crucial to understanding an infectious disease’s potential spread and informing public health interventions[3]. Analyzing $R_{0}$ can give insights into the future dynamics of a particular disease based on the parameters, subpopulations, and constants that make up any given $R_{0}$ value. Researchers calculate $R_{0}$ in models to assess transmission dynamics, compare scenarios, and evaluate the effectiveness of control measures such as vaccination or social behavior [3]. Mathematical treatments of $R_{0}$ via the ”next-generation matrix method” (NGM) have been outlined by Diekmann [2, 4, 5] and Van den Driessche [6]. Although other approaches exist for calculating $R_{0}$, the next-generation method is one of the most commonly used for ODE models[3] and so we focus on it here.
@@
Petri nets provide a promising alternative to standard Ordinary Differential Equation models in epidemiology [7, 8, 9]. Petri net models are easily visualized as causal network diagrams, and their modular structure is possible to scale to large spatial models [10]. They can also integrate continuous time models of disease spread with discrete-time events such as policy changes or interventions. Numerous studies have explored the use of Petri nets to represent and analyze SIR models, highlighting their potential for modular and large scale modeling of disease dynamics [11, 12, 13, 14, 15, 16]. However, unlike the well-developed and widely applicable techniques for computing the basic reproduction number $R_{0}$ in ODE models, no comparably general method exists for deriving $R_{0}$ directly from Petri net representations. The inability to directly calculate $R_{0}$ for Petri nets limits the analytical utility of Petri nets and poses a significant barrier to the broader adoption for the modular construction of large scale, mechanistic models of infectious disease spread.
@@
Applying the next-generation matrix method of calculating $R_{0}$ within the framework of Petri nets requires insights on how to bridge the gap between mechanistic model structure and formal spectral analysis. The contribution of this paper is introducing those insights in how the transition and transmission matrices are defined and characterizing $R_{0}$ in the context of Petri nets. Among the various formulations of the next-generation matrix method, the computational approach introduced by Diekmann et al. [2, 4, 5] is uniquely well-suited for implementation in Petri nets, as it explicitly constructs the next-generation matrix from compartmental transitions, which naturally aligns with the token flow and transition structure of Petri nets. This compatibility allows for a direct mapping from Petri net dynamics to the linearized infection and transition matrices required for $R_{0}$ calculation.
@@
In this work, we demonstrate the applicability and generality of this approach across a range of epidemiological models, including classical structures like the SIRS and SEEIR models, a compartmental COVID-19 model with various infected compartments, and a model with a nonlinear infection term. These examples collectively illustrate how the Diekmann formulation adapted to Petri nets enables rigorous and automated computation of $R_{0}$ in Petri nets, supporting their use across a broader spectrum of infectious disease modeling.
The remainder of the paper is structured as follows. In Section 2, we introduce the Petri net formalism and two ways of using PNs to implement Susceptible-Infectious-Recovered (SIR) models. One approach uses deterministic, discrete-time PNs with variable arc weights and the other uses stochastic, continuous-time PNs with constant arc weights representing the mass action of interactions among individuals in the compartments. Section 3 reviews Diekmann’s approach to calculating $R_{0}$ and shows how it applies to a range of example model types, including non-linear models and spatial patch systems. Section 4 presents a numerical verification of our approach by comparing the calculated PN $R_{0}$ values to the corresponding ODE $R_{0}$ values.
@@ 2 A Brief Overview of Petri Nets
A Petri net graph, or Petri net structure, is a weighted bipartite graph [17] defined as $n$-tuple $(P,T,A,w,x)$ where:
@@
  • $A$ is the set of arcs (edges) from places to transitions and from transitions to places in the graph $A\subset(P\times T)\cup(T\times P)$.
  • $M_{0}$ is the initial state (also known as marking), $M_{0}=[m_{1},m_{2},\ldots,m_{n}]$, where $m_{i}$ is the number of tokens in place $p_{i}$
  • $w$ : A → {1, 2, 3, . . .} is the weight function on the arcs.
  • $x$ is a marking of the set of places P; $x=[x(p_{1}),x(p_{2}),...,x(p_{n})$] $\in N^{n}$ is the row vector associated with $x$.
Tokens are assigned to places, with the initial assignment being the initial marking. The number of tokens assigned to a place is an arbitrary non-negative integer but does not necessarily have an upper bound. A transition $t_{j}\in T$ in a Petri net is said to be enabled if $x(p_{i})\geq w(p_{i},t_{j})$ for all $p_{i}\in I(t_{j})$, where $I(t_{j})$ is the set of input arcs from places to $t_{j}$. This allows us to define the state transition function, $f:{\mathbb{N}}^{n}\times T\to{\mathbb{N}}^{n}$, of Petri net $(P,T,A,w,x)$ is defined for transition $t_{j}\in T$ if and only if $x(p_{i})\geq w(p_{i},t_{j})$ for all $p_{i}\in I(t_{j})$. If $f(x,t_{j})$ is defined, then we set $x^{\prime}=f(x,t_{j})$, where $x^{\prime}(p_{i})=x(p_{i})-w(p_{i},t_{j})+w(t_{j},p_{i}),i=1,...,n$. In simple terms, a transition is enabled if the number of tokens in all places connected to that transition via an incoming arc is greater than or equal to the arc weight for the respective arc connected to the transition.
An important point to note is that there may be multiple possible PN representations for a given ODE model because there are choices in how to implement the PN model. The most common mapping method follows a method of mapping ODEs to Stochastic Petri nets (SPNs) as laid out by Baez and Biamonte [18]. An alternative method is to map the ODE to a Variable Arc Weight Petri net (VAPN) [7]. Variable arc weights Petri nets can have arc weights that depend on parameters, the number of tokens in places at specific times, and time. This method is deterministic and invertible, allowing for potentially lower computation in parameter fitting and simulations, but it is novel and not yet widely used.
@@ 3 Next-Generation Matrix For Petri Nets
We first review the next-generation method for ODEs before showing how we can apply it to PNs.
@@ 3.1 Next-Generation Matrix for Ordinary Differential Equations
The method for calculating $R_{0}$ via the Next-Generation Matrix (NGM) laid out by Diekmann et al. [2, 4, 5] defines $R_{0}$ as the spectral radius of the next-generation matrix. Specifically, two matrices (transmission and transition matrices) are defined to find $R_{0}$, starting with $T$, which is defined as the transmission matrix. The transmission matrix outlines the production of new infections. Where elements within the matrices describe how people move from non-infected compartments to infected compartments. Next, they define $\Sigma$ as the transition matrix, which describes state changes due to individuals moving to other infected compartments, dying, or becoming immune. To then find $R_{0}$ we take the dominant eigenvalue of $(-T\Sigma^{-1})$ yielding $R_{0}=\varphi(-T\Sigma^{-1})$.
Alternatively, van den Driessche and Watmough [6] outline a similar definition of $R_{0}$. Let $x=(x_{1},…,x_{n})^{t}$, with each $x_{i}\geq 0$, be the number of individuals in each compartment. Let $n$ be the number of compartments with the first $m$ compartments being the infected compartments. We define $X_{s}$ as the set of all disease free states such that $X_{s}=\{x\geq 0|x_{i}=0,i=1,...,m\}$. Then we define the transmission matrix, matrix $\mathcal{F}_{i}(x)$ to be the rate of appearance of new infections in compartment $i$. We use two matrices $\mathcal{V}^{+}_{i}(x)$ and $\mathcal{V}^{-}_{i}(x)$ to describe the rate of transfer of individuals into compartment $i$ by all other means and the rate of transfer of individuals out of compartment $i$, respectively. From $\mathcal{V}^{+}_{i}(x)$ and $\mathcal{V}^{-}_{i}(x)$, we get the transition matrix, $\mathcal{V}_{i}=\mathcal{V}^{-}_{i}(x)-\mathcal{V}^{+}_{i}(x).$ The disease model itself has non-negative initial conditions with the system of equations following
$\dot{x_{i}}$ $=$ $f_{i}(x)=\mathcal{F}_{i}(x)-\mathcal{V}_{i}(x),i=1,...n.$ (1)
While some of these assumptions are trivial when modeled in Petri nets, others, like derivations, need to be inspected more closely. For reference, we state the five mathematical assumptions here exactly as they are given in van den Driessche and Watmough [6] followed by brief comments. We will cover these assumptions for Petri nets in the following Section 3.
@@
    If $x\geq 0$, then $\mathcal{F}_{i},\mathcal{V}^{+}_{i},\mathcal{V}^{-}_{i}\geq 0$ for $i=1,…,n$. Each subsequent function $\dot{x_{i}}$ represents a transfer of individuals, thus they are all non negative.
@@
    If $x_{i}=0$ then $\mathcal{V}^{-}_{i}=0$. In particular, if $x\in X_{s}$ , then $\mathcal{V}^{-}_{i}=0$ for $i=1,…,m$. If a compartment is empty, then there can be no transfer of individuals out of the compartment.
@@
    If $x\in X_{s}$ then $\mathcal{F}_{i}(x)=0$ and $\mathcal{V}^{+}_{i}(x)=0$ for $i=1,…,m$. This simply states that if the population is disease-free, it stays disease-free. Meaning no, density-independent infections population immigrates in. This becomes important when dealing with patch models, as seen in Example 4.6.1.
@@
    If $\mathcal{F}(x)$ is set to zero, then all eigenvalues of $Df(x_{0})$ have negative real parts, where $Df(x_{0})$ is the derivative $[\frac{\partial f_{i}}{\partial x_{j}}]$ evaluated at the disease free equilibrium (DFE), $x_{0}$.
@@
These assumptions allow for the proof of Lemma 1 from van den Driessche and Watmough [6] and using Lemma 1, van den Driessche and Watmough go on to prove $\varphi(FV^{-1})=R_{0}$.
3.2 Next-Generation Matrix for Petri Nets
Since many of the papers we draw on for examples below use the van den Driessche and Watmough notation, we will lay out the Next-Generation Matrix for Petri nets (NGMPN) following similar notation. To help in visualizing some of the concepts outlined in this process Figure 1 will be used as a basic VAPN model with non-infected places ($N_{1},N_{2}$) and infected places ($I_{1},I_{2},I_{3}$). Examples using SPNs will be given in the Case Studies section under various cases.

### v3

# petri nets; reproduction number; SIR model; disease spread; next-generation method; ordinary differential equations; epidemiology; modeling

@@ Abstract
The basic reproduction number ($R_{0}$) is an epidemiological metric that represents the average number of new infections caused by a single infectious individual in a completely susceptible population. The methodology for calculating this metric is well-defined for numerous model types, including, most prominently, Ordinary Differential Equations (ODEs). The basic reproduction number is used in disease modeling to predict the potential of an outbreak and the transmissibility of a disease, as well as by governments to inform public health interventions and resource allocation for controlling the spread of diseases. A Petri Net (PN) is a directed bipartite graph where places, transitions, arcs, and the firing of the arcs determine the dynamic behavior of the system. Petri Net models have been an increasingly used tool within the epidemiology community. However, no generalized method for calculating $R_{0}$ directly from PN models has been established. Thus, in this paper, we establish a generalized computational framework for calculating $R_{0}$ directly from Petri Net models. We adapt the next-generation matrix method to be compatible with multiple Petri Net formalisms, including both deterministic Variable Arc Weight Petri Nets (VAPNs) and stochastic continuous-time Petri Nets (SPNs). We demonstrate the method’s versatility on a range of complex epidemiological models, including those with multiple strains, asymptomatic states, and nonlinear dynamics. Crucially, we numerically validate our framework by demonstrating that the analytically derived $R_{0}$ values are in strong agreement with those estimated from simulation data, thereby confirming the method’s accuracy and practical utility.
keywords:
petri nets; reproduction number; SIR model; disease spread; next-generation method; ordinary differential equations; epidemiology; modeling
[…]
@@ 1 Introduction
Petri Nets are a promising modeling framework for infectious disease systems that are difficult for standard Ordinary Differential Equation (ODE) models Reckell et al. (2024); Segovia (2025); Peleg et al. (2005). Petri Net models are easily visualized as causal network diagrams, and their modular structure can be scaled to large spatial models Chiaradonna et al. (2024). They can also integrate continuous-time models of disease spread with discrete-time events such as policy changes or interventions. Numerous studies have explored the use of Petri Nets to represent and analyze SIR models, highlighting their potential for modular and large-scale modeling of disease dynamics Bahi-Jaber and Pontier (2003); Peng et al. (2021); Kong et al. (2021); Connolly et al. (2022); Libkind et al. (2022); Segovia (2025).
However, Petri Nets have lacked a generalized method for deriving the basic reproduction number, $R_{0}$. In epidemiology, the basic reproduction number ($R_{0}$) refers to the expected number of new cases that one infected individual will cause in a population where the entire population is susceptible Dietz (1993); Diekmann et al. (1990). As such, $R_{0}$ is a fundamental concept in mathematical epidemiology for all communicable diseases. While the Next Generation Matrix (NGM) framework is standard for ODEs, its adaptation to Petri Nets is not yet widely established. Concurrent with our work Reckell et al. (2025), Segovia has also proposed a formal adaptation of the next-generation matrix method to Petri Nets, focusing on a geometric interpretation of the matrix entries Segovia (2025). Our paper further advances this emerging topic by establishing a universal, computationally-oriented framework that offers distinct contributions. Compared to the geometric interpretation Segovia (2025), our framework provides a higher degree of formalization for complex models through a direct method, a wider range of applications, and a more comprehensive verification method. Specifically, our approach is demonstrated on both traditional Stochastic Petri Nets (SPNs) and, crucially, the more flexible Variable Arc Weight Petri Nets (VAPNs). This VAPN compatibility (which can be applied to a wider range of models, such as those in Section 4.6) enables our framework to compactly model complex, nonlinear dynamics that are challenging for path-based methods. Furthermore, we provide, to our knowledge, the first robust numerical verification of an NGM Petri Net framework (Section 5), confirming its accuracy and practical utility.
The inability to directly calculate $R_{0}$ for Petri Nets limits their analytical utility and poses a significant barrier to the broader adoption of the modular construction of large-scale, mechanistic models of infectious disease spread. An accurate calculation of $R_{0}$ is therefore crucial for understanding the potential spread of an infectious disease and informing public health interventions Delamater et al. (2019). Analyzing $R_{0}$ can give insights into the future dynamics of a particular disease based on how its value depends on the specific parameters, subpopulations, and constants that characterize the system at hand. Researchers calculate $R_{0}$ in models to assess transmission dynamics, compare scenarios, and evaluate the effectiveness of control measures such as vaccination or social behavior Delamater et al. (2019).
Applying the NGM method for calculating $R_{0}$ within the framework of Petri Nets requires insights bridging the gap between mechanistic model structure and formal spectral analysis. Mathematical treatments of $R_{0}$ via the next-generation matrix method have been outlined by Diekmann Diekmann et al. (1990); Diekmann and Heesterbeek (2000); Diekmann et al. (2010) and Van den Driessche Van den Driessche and Watmough (2002). Although other approaches exist for calculating $R_{0}$, such as stability analysis of the disease-free equilibrium (DFE) Van den Driessche and Watmough (2002), the direct definition from survival functions Hethcote (2000), and estimation from final size relations Brauer et al. (2008), the next-generation method is one of the most commonly used for ODE models Delamater et al. (2019).
@@
In this paper, we show how to define and characterize the transition and transmission matrices necessary to calculate $R_{0}$ in Petri Nets. Among the various formulations of the NGM method, the computational approach introduced by Diekmann et al. Diekmann et al. (1990); Diekmann and Heesterbeek (2000); Diekmann et al. (2010) is uniquely well-suited for implementation in Petri Nets, as it explicitly constructs the next-generation matrix from compartmental transitions, which naturally aligns with the token flow and transition structure of Petri Nets. This compatibility enables a direct mapping from Petri Net dynamics to the linearized infection and transition matrices necessary for calculating $R_{0}$.
@@
We demonstrate the applicability and generality of this approach across a range of epidemiological models, including classical structures such as the SIRS, SEEIR models, a compartmental COVID-19 model with various infected compartments, and a model featuring a nonlinear infection term. These examples collectively illustrate how the Diekmann formulation, adapted to Petri Nets, enables the rigorous and automated computation of $R_{0}$ in Petri Nets, supporting their use across a broader spectrum of infectious disease modeling.
A premise of this work is that epidemiological models originally formulated as systems of ODEs can be systematically represented as Petri Nets, allowing for a more modular and visual approach to modeling. There are two main methodologies for mapping from PNs to ODEs that facilitate this translation. The first, commonly used for Stochastic Petri Nets (SPNs), maps the system based on mass-action principles. In this approach, outlined by Baez and Biamonte Baez and Biamonte (2012), the state variables (compartments) of the ODE system correspond to the places in the PN, and each term in the differential equations, representing an infection or recovery event, corresponds to a transition governed by a stochastic rate. The second method maps ODEs to a deterministic Variable Arc Weight Petri Net (VAPN). Here, the entire mathematical expression for the rate of change between two compartments can be directly encoded as a functional arc weight on the arc connecting the corresponding places. Both the SPN and VAPN have been numerically verified for the SIRS system Reckell et al. (2024). Our framework is intentionally designed to be compatible with both of these ODE to PN mapping formalisms. By demonstrating that our method for determining $R_{0}$ for Petri Nets yields equivalent $R_{0}$ values for PN models derived from either method, we ensure its broad applicability and relevance to the wider epidemiological modeling community.
The remainder of the paper is structured as follows. In Section 2, we introduce the Petri Net formalism and discuss two approaches for using PNs to implement Susceptible–Infectious–Recovered (SIR) models. One approach uses deterministic, discrete-time PNs with variable arc weights, and the other uses stochastic, continuous-time PNs with constant arc weights representing the mass action of interactions among individuals in the compartments. Section 3 presents our NGM framework for Petri Nets. In Section 4, we demonstrate the broad applicability of this framework through a series of case studies on diverse and complex epidemiological models. Finally, a key contribution of this work is presented in Section 5, which provides a robust numerical verification of our framework. By simulating the Petri Net models directly and applying established statistical methods to the output data, we demonstrate a strong concordance between our analytically derived $R_{0}$ and the epidemiologically observed $R_{0}$ from simulations, thereby validating the accuracy of our approach.
@@ 2 A Brief Overview of Petri Nets
A Petri Net graph, or Petri Net structure, is a weighted bipartite graph Cassandras and Lafortune (2008) defined as an $n$-tuple $(P,T,A,w,x)$ where:
@@
  • AA is the set of arcs (edges) from places to transitions and from transitions to places in the graph $A\subset(P\times T)\cup(T\times P)$.
  • $w$ : $A$ → {1, 2, 3, …} is the weight function on the arcs.
  • $x$ is a marking of the set of places $P$; $x_{i}=[x(p_{1}),x(p_{2}),\ldots,x(p_{n})$] $\in{\mathbb{N}}^{n}$ is the row vector associated with $x$. The term $x_{i}$ refers to marking of $x$ at state $i$. Note that $x_{0}$ is the initial state (also known as marking).
Tokens are assigned to places, with the initial assignment being the initial marking. The number of tokens assigned to a place is an arbitrary non-negative integer, but it does not necessarily have an upper bound. A transition $t_{j}\in T$ in a Petri Net is said to be enabled if $x(p_{i})\geq w(p_{i},t_{j})$ for all $p_{i}\in I(t_{j})$, where $I(t_{j})$ is the set of input arcs from places to $t_{j}$. This allows us to define the state transition function, $f:{\mathbb{N}}^{n}\times T\to{\mathbb{N}}^{n}$, of Petri Net $(P,T,A,w,x)$ is defined for transition $t_{j}\in T$ if and only if $x(p_{i})\geq w(p_{i},t_{j})$ for all $p_{i}\in I(t_{j})$. If $f(x,t_{j})$ is defined, then we set $x^{\prime}=f(x,t_{j})$, where $x^{\prime}(p_{i})=x(p_{i})-w(p_{i},t_{j})+w(t_{j},p_{i}),i=1,\ldots,n$. In simple terms, a transition is enabled if the number of tokens in all places connected to that transition via an incoming arc is greater than or equal to the arc weight for the respective arc connected to the transition.
An important point to note is that there may be multiple possible PN representations for a given ODE model because there are choices in how to implement the PN model. The most common mapping method involves mapping ODEs to Stochastic Petri Nets (SPNs), as outlined by Baez and Biamonte Baez and Biamonte (2012). An alternative method is to map the ODE to a Variable Arc Weight Petri Net (VAPN) Reckell et al. (2024). Variable arc weight Petri Nets can have arc weights that depend on parameters, the number of tokens in places at specific times, and time. VAPNs allow for arc weights to be defined as functions that change over time, which can be linear or nonlinear. This method is deterministic and invertible, allowing for potentially lower computational requirements in parameter fitting and simulations, but it is novel and not yet widely used.
The inclusion of VAPNs is a key feature of our framework’s generality. While SPNs excel at modeling mass-action kinetics, VAPNs can represent complex, nonlinear transmission functions (such as frequency-dependent or behavior-driven rates) as single, functional arc weights. This allows for more elegant and compact models. Specifically, VAPNs can directly model mechanisms that are not based on mass-action principles, a limitation of traditional SPNs. As we will demonstrate, our method of identifying the arcs that contribute to the $\mathcal{F}$ and $\mathcal{V}$ matrices works intrinsically on these functional weights, a task not explicitly addressed by methods tailored to the constant-rate transitions of traditional SPNs.
@@ 3 Next-Generation Matrix for Petri Nets
We first review the next-generation method for ODEs before demonstrating how it can be applied to PNs.
@@ 3.1 Next-Generation Matrix for Ordinary Differential Equations
The Next-Generation Matrix (NGM) is a standard and widely-used technique in mathematical epidemiology for calculating $R_{0}$ from compartmental models. For additional methods for calculating $R_{0}$, see Martcheva (2015). The method for calculating $R_{0}$ via the NGM laid out by Diekmann et al. Diekmann et al. (1990); Diekmann and Heesterbeek (2000); Diekmann et al. (2010) defines $R_{0}$ as the spectral radius of the next-generation matrix. Specifically, two matrices (the transmission and transition matrices) are defined to calculate $R_{0}$, starting with $T$, which is defined as the transmission matrix. The transmission matrix outlines the production of new infections. Where elements within the matrices describe how people move from non-infected compartments to infected compartments. Next, they define $\Sigma$ as the transition matrix, which describes the state changes that occur when individuals move to other infected compartments, die, or become immune. To then find $R_{0}$ we take the dominant eigenvalue of $(-T\Sigma^{-1})$ yielding $R_{0}=\varphi(-T\Sigma^{-1})$.
Alternatively, van den Driessche and Watmough Van den Driessche and Watmough (2002) outline a similar definition of $R_{0}$. Let $x=(x_{1},…,x_{n})^{t}$, with each $x_{i}\geq 0$, be the number of individuals in each compartment. Let $n$ be the number of compartments, with the first $m$ compartments being the infected compartments. We define $X_{s}$ as the set of all disease-free states such that $X_{s}=\{x\geq 0|x_{i}=0,i=1,\ldots,m\}$. Then we define the transmission matrix, matrix $\mathcal{F}_{i}(x)$, to be the rate of appearance of new infections in compartment $i$. We use two matrices $\mathcal{V}^{+}_{i}(x)$ and $\mathcal{V}^{-}_{i}(x)$ to describe the rate of transfer of individuals into compartment $i$ by all other means and the rate of transfer of individuals out of compartment $i$, respectively. From $\mathcal{V}^{+}_{i}(x)$ and $\mathcal{V}^{-}_{i}(x)$, we get the transition matrix, $\mathcal{V}_{i}=\mathcal{V}^{-}_{i}(x)-\mathcal{V}^{+}_{i}(x).$ The disease model itself has non-negative initial conditions with the system of equations following
$\dot{x_{i}}$ $=f_{i}(x)=\mathcal{F}_{i}(x)-\mathcal{V}_{i}(x),i=1,{\ldots}n.$ (1)
While some of these assumptions are trivial when modeled in Petri Nets, others, such as derivations, require closer inspection. For reference, we state the five mathematical assumptions here exactly as they are given in van den Driessche and Watmough Van den Driessche and Watmough (2002), followed by brief comments. We will cover these assumptions for Petri Nets in the following Section 3.2.
@@
    If $x\geq 0$, then $\mathcal{F}_{i},\mathcal{V}^{+}_{i},\mathcal{V}^{-}_{i}\geq 0$ for i=1,,ni=1,…,n. Each subsequent function $\dot{x_{i}}$ represents a transfer of individuals, thus they are all non-negative.
@@
    If $x_{i}=0$ then $\mathcal{V}^{-}_{i}=0$. In particular, if $x\in X_{s}$, then $\mathcal{V}^{-}_{i}=0$ for i=1,,mi=1,…,m. If a compartment is empty, then there can be no transfer of individuals out of the compartment.
@@
    If $x\in X_{s}$ then $\mathcal{F}_{i}(x)=0$ and $\mathcal{V}^{+}_{i}(x)=0$ for i=1,,mi=1,…,m. This states that if the population is disease-free, it stays disease-free. Meaning no density-independent infectious population immigrates in. This becomes important when dealing with patch models, as seen in Section 4.7.
@@
    If $\mathcal{F}(x)$ is set to zero, then all eigenvalues of $Df(x_{0})$ have negative real parts, where $Df(x_{0})$ is the derivative $\frac{\partial f_{i}}{\partial x_{j}}]$ evaluated at the disease free equilibrium (DFE), $x_{0}$.
@@
These assumptions enable the proof of Lemma 1 from van den Driessche and Watmough Van den Driessche and Watmough (2002). Using Lemma 1, van den Driessche and Watmough subsequently prove that $\varphi(FV^{-1})=R_{0}$.
3.2 Constructing the Next-Generation Matrix for Petri Nets
Since many of the papers we draw on for examples below use the van den Driessche and Watmough notation, we will lay out the Next-Generation Matrix for Petri Nets (NGMPN) following similar notation. To help visualize some of the concepts outlined in this process, Figure 1 will be used as a basic VAPN model with non-infected places ($N_{1},N_{2}$) and infected places ($I_{1},I_{2},I_{3}$). Examples using SPNs will be provided in the Case Studies section under in the first two model examples. This process of separating disease transmission dynamics from other state transitions within the infected population is conceptually similar to the framework of ‘Kermack–McKendrick module’ proposed by Segovia (2025). In that work, the Petri Net is partitioned into ‘susceptible’, ‘infection-process’, and ‘infection’ modules, where the $F$ matrix is derived from the infection-process module and the $V$ matrix is derived from the infection module. Our approach implements this separation by identifying specific arcs and transitions corresponding to new infections ($\mathcal{F}_{i}(x)$) versus those representing recovery, death, or movement between infected states ($\mathcal{V}_{i}(x)$). It is important to note that, as with any $R_{0}$ calculation in any modeling formalism (including ODEs), the modeler must first define the set of infected places or compartments. This is a necessary prerequisite of the epidemiological modeling process itself. Our framework’s “intuitiveness” lies in how, given this standard partition of places, the $F$ and $V$ matrices can be directly and systematically constructed by identifying the specific arcs that cross this partition, as we will formally define.